OpenStudy (anonymous):

Find the roots of : x^2+11x=-30 3x^2+2x-4=0 5 sqrt x -x=6 x^2-8=17 CAN ANYBODY HELP ME PLEASE WORK THESE OUT !?!

OpenStudy (anonymous):

This site explains how to find roots and gives some good examples: http://www.freemathhelp.com/finding-roots.html

OpenStudy (madhatter):

1st one is x=60

OpenStudy (anonymous):

Hmm, I am not getting 60, perhaps if you show your work how you came about that @madhatter ? When I do it I get a simplified equation of: x^2+11x=-30 --> (x+5)(x+6) meaning its roots are at x=-5 and x=-6

OpenStudy (madhatter):

2nd one is x=24

OpenStudy (madhatter):

just do it like a norm equation its easy

OpenStudy (anonymous):

@madhatter If you plug 24 into the second equation... 3x^2+2x-4=0 3(24)^2+2(24)-4=0 1728+48-4=0 This is not right :/ as it does not equal 0, might be useful if you show your work for how you are getting those answers for @Rea731

OpenStudy (anonymous):

@edlevin the first problem you did is correct

OpenStudy (anonymous):

Yup, if you follow the same procedure for the others you should find those values also, if you would like to post what you did I can check them for you :)

OpenStudy (anonymous):

I am having a problem solving the second one